OpenStudy (anonymous):

could you please help, make c the subject of A= 1/2(b+c)h

OpenStudy (anonymous):

\[\Large \begin{array}{l} A = \frac{1}{2}(b + c)h\\ c = \frac{{2A}}{h} - b \end{array}\]

OpenStudy (anonymous):

my friend thinks the answer is c=2A-bh/h

OpenStudy (across):

Sadly, your friend is incorrect!

OpenStudy (anonymous):

\[\Large \begin{array}{l} A = \frac{1}{2}(b + c)h\\ (b + c) = \frac{A}{{\frac{1}{2}h}}\\ c = \frac{{2A}}{h} - b \end{array}\] Let's assign some arbitrary values and test it: b = 2, c = 3, h = 4. A = (1/2)(2+3)4 = 10 c = ((2*10)/4) - 2 = (20/4) - 2 = 5 - 2 = 3. Seems legit.

OpenStudy (anonymous):

easy: 1. mutiply both sides by 2, 2. divide boht sides by h 3. subtract b from both sides done!

OpenStudy (anonymous):

could someone please guide me through each step